. A car of mass M traveling with velocity v strikes a car of mass M that is at rest. The two cars’ bodies mesh in the collision. The loss of the kinetic energy the moving car undergo in the collision is


1
Expert's answer
2020-08-06T16:44:50-0400

Using the momentum conservation law, we can write:


"Mv + 0 = (M+M)V = 2MV"

where "Mv" and momentums of the first and second car before the collision respectively, and "V" is a speed of the joined cars after the collision. Expressing "V", obtain:


"V = \dfrac{v}{2}"

The initial kinetic energy of the first car is:


"K_1 = \dfrac{Mv^2}{2}"

The final kinetic energy of the first car is:

"K_2 = \dfrac{MV^2}{2} = \dfrac{Mv^2}{2\cdot 4} = \dfrac{Mv^2}{8}"

The loss of the kinetic energy is:


"1-\dfrac{K_2}{K_1} =1- \dfrac14 = \dfrac34"

Thus, 3/4 of the initial kinetic energy was lost in the collision.


Answer. 3/4.